twice the sum of 4 and 3 times some number is 34. what is the number? I need this in step process so I will understand how to do it please
Answer by rapaljer(4671)

I'm not sure my steps are the same as your steps, but you can adapt what I do.

Step 1: Let x = the number

Step 2: Write the equation:
"3 times some number" means 3x

"the sum of 4 and 3 times some number" means
3x+4

"twice the sum of 4 and 3 times some number means
2(3x+4)

This equals 34
2(3x+4) = 34
6x +8=34

Subtract 8 from each side to "undo" +8:
6x+8-8= 34-8
6x=26

Divide by 6:
x= 26/6=13/3

Step 4: Answer the question.
The number is 13/3.

Step 5: Check.
3 times the number is 3*13/3, which is 13.

Four more than this is 17.

Twice this is 34, so it does check.

For additional explanation on Equation Solving, please see my own website. You can find my website by doing a Google search for my name "Dr. Robert J. Rapalje". Look for "Rapalje Homepage", which should be the first item on the search. Once you find my homepage, click on "Basic, Intermediate, and College Algebra: One Step at a Time." Look in Basic Algebra in Chapter 1 for Section 1.08 Equation Solving. There you will find a detailed explanation, and solutions to problems in LIVING COLOR!!

I have many other topics explained in this way from Basic, Intermediate, and College Algebra. Please check out my website to see what I have. It's all FREE, and if you send me an Email, I'll be glad to help you find something. I'm not selling anything. If you get a good explanation, math is usually not as hard as you think it is.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus